# Po-Heng Chou1, Chiapin Wang2, Kuan-Hao Chen2, and Wei-Chen Hsiao2

### v1

@@ Abstract
In this paper, we propose a reinforcement learning based beam weighting framework that couples a policy network with an augmented weighted least squares (WLS) estimator for accurate and low-complexity positioning in multi-beam LEO constellations. Unlike conventional geometry or CSI-dependent approaches, the policy learns directly from uplink pilot responses and geometry features, enabling robust localization without explicit CSI estimation. An augmented WLS jointly estimates position and receiver clock bias, improving numerical stability under dynamic beam geometry. Across representative scenarios, the proposed method reduces the mean positioning error by 99.3% compared with the geometry-based baseline, achieving 0.395 m RMSE with near real-time inference.
@@ I Introduction
Beyond the physical layer, system-level frameworks have also been explored. A symbiotic radio approach using collaborative DRL has been proposed for intelligent resource optimization in non-terrestrial networks (NTNs) [3]. Standardization and system-level research on integrated terrestrial and non-terrestrial networks (TN-NTN) localization have highlighted key challenges in synchronization, interference, and robustness [1]. Moreover, federated learning (FL) frameworks have been investigated for NTNs to enhance scalability and privacy in multi-tier deployments [9].
Despite these advances, most existing solutions still rely heavily on explicit CSI or involve computationally intensive optimization, limiting real-time applicability for resource-constrained LEO platforms. Traditional statistical estimation methods, such as weighted least squares (WLS) [10], widely used in positioning problems, provide reliable accuracy but lack adaptability in dynamic multi-beam environments. Meanwhile, DRL has proven effective in sequential decision-making tasks [11], yet its integration with lightweight positioning remains underexplored in NTNs.
@@ IV-C Training Procedure
where $\delta_{t}=y_{t}-Q_{\theta}(\bm{s}_{t},\bm{w}_{t})$. Gradients are clipped with $\|\nabla_{\theta}\|\leq 5$ to prevent divergence, and $\epsilon$ decays exponentially from 1.0 to 0.01 during training. Convergence is typically achieved within 800 to 1000 episodes, after which the mean positioning error stabilizes below 0.05 m. Furthermore, the learned beam weights exhibit strong interpretability: high-SINR beams receive larger weights, while redundant or correlated beams are automatically suppressed, indicating that the agent implicitly learns the beam-geometry reliability without explicit CSI modeling.

### v2

@@ Abstract
In this paper, we propose a deep reinforcement learning (DRL)-based beam weighting framework that couples a policy network with an augmented weighted least squares (WLS) estimator for accurate and low-complexity positioning in multi-beam LEO constellations. Unlike conventional geometry or channel state information (CSI)-dependent approaches, the policy learns directly from uplink pilot responses and geometry features, enabling robust localization without explicit CSI estimation. An augmented WLS jointly estimates position and receiver clock bias, improving numerical stability under dynamic beam geometry. Across representative scenarios, the proposed method reduces the mean positioning error by 99.3% compared with the geometry-based baseline, achieving 0.395 m RMSE with near real-time inference.
@@ I Introduction
Beyond the physical layer, system-level frameworks have also been explored. A symbiotic radio approach using collaborative DRL has been proposed for intelligent resource optimization in non-terrestrial networks (NTNs) [3]. Standardization and system-level research on integrated terrestrial and non-terrestrial networks (TN-NTN) localization have highlighted key challenges in synchronization, interference, and robustness [1]. Moreover, federated learning frameworks have been investigated for NTNs to enhance scalability and privacy in multi-tier deployments [9].
Despite these advances, most existing solutions still rely heavily on explicit channel state information (CSI) or involve computationally intensive optimization, limiting real-time applicability for resource-constrained LEO platforms. Traditional statistical estimation methods, such as weighted least squares (WLS) [10], widely used in positioning problems, provide reliable accuracy but lack adaptability in dynamic multi-beam environments. Meanwhile, DRL has proven effective in sequential decision-making tasks [11], yet its integration with lightweight positioning remains underexplored in NTNs.
@@ IV-C Training Procedure
where $\delta_{t}=y_{t}-Q_{\theta}(\bm{s}_{t},\bm{w}_{t})$. Gradients are clipped with $\|\nabla_{\theta}\|\leq 5$ to prevent divergence, and $\epsilon$ decays exponentially from 1.0 to 0.01 during training. Convergence is typically achieved within 800 to 1000 episodes, after which the mean positioning error stabilizes below 0.5 m. Furthermore, the learned beam weights exhibit strong interpretability: high-SINR beams receive larger weights, while redundant or correlated beams are automatically suppressed, indicating that the agent implicitly learns the beam-geometry reliability without explicit CSI modeling.